FEL Upgrade Project Weekly Brief
November 13 – 17, 2000


We complete a successful Fall user run at the end of today’s second shift.  Thanks to all of the FEL team and our Fall users for their dedicated work.  Good progress was made this week on incorporating the configuration changes that were discussed at the October 23 review of the FEL Upgrade design.

NOTE:  Due to the Thanksgiving Holidays (Nov. 23-24) there will be no FEL Project weekly brief next week


The cost performance data for the month of October was collected from our Finance Dept. and reviewed with the WBS Managers in preparation for writing the October Monthly Report.  Presentations on the FEL Project were made at this week’s meeting of SURA’s Board of Trustees held in Washington, DC.

WBS 3 (Beam Physics)

Analysis of injector simulations at 135 pC has led to a recommendation for the use of sector-dipole injection line geometry in the upgrade. This will be incorporated in the Revision 1.1 release of the machine design.

Design of Revision 1.1 has started; this will feature the use of sector dipoles in the injection/extraction line, space for use of 7-cell modules everywhere, revised magnet field roll-off values, and will accommodate the footprint of the adopted 32 m near-concentric optical cavity resonator.  Initial studies of the injection line geometry for this design has provided an explanation for the reinjection steering observed in the IR Demo - it is due to the use of sector dipoles in the reinjection line - and suggest this will be less significant in the Upgrade, where the bend angles are smaller.  Beam optics with revised linac warm region slot lengths (allowing use of
7-cell modules) has been developed; integration with the recirculator design is underway. A plot of beam envelopes in included below.

WBS 4 (Injector)

The second nitrogen implanted electrode was run at 30 MV/m in the field emission test system. Its base line current was around 180 pA which was higher than the first (55 pA) but still well below the limits of acceptability.  The electrode was run for 7 hours at 30 MV/m showing a higher but less active baseline current.


Visited the waveguide supplier for kick off of the contract.  Vendor looks capable and well suited to supply parts.  First Article helium vessels shipped today from PHPK, vessel vendor.

WBS 6 (RF)

Continued installation and check-out activities in Zones 3 and 4

WBS 8 (I&C)

The optical beam position monitoring system (starting with can 6) is now operational. There was a fair amount of trouble getting the 32 channel VME sample and hold board to work in the "real" crate. This new design Was
tested in a crate without a functioning IOC and worked fine, when it was inserted into the crate in zone 2, the IOC failed to boot properly. The problem was tracked down to the OMS stepper motor controller - if a blank card extender was placed to the left of the OMS card bus noise would couple in to the card and cause a failure. The cards were rearranged and all worked fine. The quad pyro detector electronics perform extremely well – the balanced signal is > 5 volts with rise times of ~ 15 microseconds. The control screen displays both the position and incident power

WBS 9 (Transport)

o Six vendors have the preliminary dipole package for budgetary quotes as well as review and comment.  George Biallas walked one vendor through the package during the same trip as a Berkeley Lab magnet review.  No comments thus far.
 o GW (optical cavity dipole) prototype drawings are continuing to be developed.  3D magnetic models of the GW have been run using three different materials for the Purcell Gap and at multiple field strengths.  K1 and effective lengths have been calculated for all cases.
o Next Priority:
o Determine effective length and K1 from existing baseline-GX model
o Build brand new 3D model of modified-GX accommodating UV branching QX (3.125" Quad) Magnets
o Prototype Fabrication: vendor has started machining the four steel pole tips with a projected delivery date of November 27th.
o Budgetary Quote: An encouraging quote was received from Master Machine who proposed a mix of wire EDM and CNC machining.  Five additional requests are still outstanding.
o Measurement Probe: The litz wire coils are being built in-house by a technician in the HPEE Group.  Printed circuit boards for these coils were made by Kim Ryan and Ty Pate. Drawings are progressing and a meeting was held to discuss design, machining, and assembly issues.

WBS 10 (Wiggler)

GS (dispersion dipole for optical klystron) Vendor has received conductor.  The steel was sent out yesterday.  We are writing up a description of the magnet measurements to send STI Optronics for a cost estimate.

WBS 11 (Optics)

The drawings for the mirror test stand were signed off and are being sent out for bids.  All the optical windows for the test stand have been ordered.  Some of the ancillary equipment, such as another vacuum pump, is
also being ordered.  Our attention now turns to completing the procurements for the mirror test stand, and integration of the mounted mirrors into the optical cavity vacuum vessel.  We've met with our designer and he's begun work on the conceptual layout.  We've also had a telephone conference with the engineer at AES responsible for finalizing the design for the HR mirror.

We received the quad pyros and associated electronics from the I & C group, (c.f. WBS 8 for more details) and this was installed onto the O-BPM mechanical assembly in the optics control room (OCR). It has resulted
in a much improved ability to determine and fix the beam position between the OCR and ModeMaster at the end-of-line (EOL).  During the next two weeks we will add the electronics to the O-BPM assembly that exists in the accelerator vault, and in the EOL diagnostics hutch.

We installed the pyro array, with a magnifying optic to increase the resolution, on the spectrograph port, in order to monitor the FEL spectrum in real time.  This has been only marginally useful, and needs the addition
of better triggering and readout electronics to take full advantage of it.

We received the custom-built enclosure for the OCR optical table, and it will be installed soon.  It will make purging the table far easier and more effective.


This was the last week of FEL operations for the Fall User run.  The FEL ran well and very stably all week. Operations were devoted to providing beam to G. Luepke (College of William and Mary) to study the dynamics of H-implanted defects in Si.  It pushes the performance of the IR Demo to the limits in terms of amplitude, wavelength, and position stability.  So far, we've been able to meet these requirements.

Privacy and Security Notice